-1/8(y+3)=-1/5(y+1/6)

Answer by unlockmath(1688)

Hello,
Let's multiply this -1/8(y+3)=-1/5(y+1/6) by -40 to get:
5(y+3)=8(y+1/6)
Expand:
5y+15=8y+4/3
Multiply be 3:
15y+45=24y+4
Subtract 4 and 15y to get:
41=9y
Divide by 9:
y=41/9
Make sense?
RJ
www.math-unlock.com